# FIITED: Fine-grained embedding dimension optimization during training for recommender systems

### v1

@@ 5.1 Experimental setup
Figure 8: Average training/inference time per iteration of different models on Criteo Kaggle dataset (batch size=4096).
5.2 Results
5.2.1 Training memory saving
@@
Public models For the public models, the criterion for model quality is prediction accuracy. According to Zhou et al. (2018) and Song et al. (2019), an accuracy loss larger than 0.1%-level is considered significant for click-through-rate prediction tasks. Table 2 shows the maximum pruning ratio of FIITED on different datasets, compared with AdaEmbed and ESAPN. On the three click-through-rate prediction datasets (i.e., Criteo Kaggle, Avazu and Criteo Terabyte), we reported the maximum pruning ratio without significant accuracy loss, while on the MovieLens-20M dataset we reported the maximum reduction ratio to obtain a comparable accuracy to ESAPN. Note that the default embedding dimension we used for FIITED on the MovieLens-20M dataset is 1/2 the size of the original maximum embedding dimension size used in ESAPN, and our reduction ratio is calculated based on that. On all four datasets, FIITED is able to prune more embeddings than the baselines while maintaining the model quality, and manual selection of pruning ratios performs a bit better than dynamical selection. Figure 8 shows the validation accuracy comparison of FIITED, AdaEmbed and Mixed Dimension on the Criteo Kaggle dataset using different pruning ratios . FIITED consistently achieves better accuracy than the baselines, and the two pruning ratio selection approaches (manual and dynamical) yield similar results. On the Criteo Terabyte dataset, FIITED is able to achieve a modest 0.1% accuracy gain under a pruning rate of 93.75%.
Table 2: Maximum embedding size reduction ratio on different datasets. ESAPN is only tested on the MovieLens-20M dataset, because its design only considers two types of embeddings, user embeddings and item embeddings, and does not apply to the other three datasets.
Model Criteo Kaggle Avazu Criteo Terabyte MovieLens-20M
AdaEmbed 98% 92.5% 99.5% 72.5%
ESAPN - - - 51.9%
FIITED-Manual 99% 95% 99.75% 81.25%
FIITED-Dynamical 99% 93.75% 99.75% 77.5%
@@
Production models The criterion for model quality is Normalized Entropy (NE) (He et al., 2014; Qin et al., 2020) on the evaluation set; a lower NE indicates a better model. The effect of pruning is evaluated by NE loss, i.e., the percentage change in NE after model pruning compared to an unpruned model. A positive NE loss indicates worsened model quality, and an NE loss bigger than 0.02% is generally considered significant. The results are shown in Figure 9. For both Model A and Model B, FIITED can achieve better NE than AdaEmbed given the same average pruning ratio. For Model A, without incurring significant NE loss ($>$0.02%), FIITED can prune 65%-75% of the embeddings, while AdaEmbed can only prune 50%-65%. For Model B, FIITED can prune up to 50% of the embeddings, while AdaEmbed can only prune 30%-40% of the embeddings without causing significant NE loss.
@@
For Model A, we also compare two different dimension pruning ratio settings with the same average pruning ratio. At 75% average pruning ratio, the “steeper” setting achieved a better NE (+0.02% VS. +0.03%) than the “flatter” setting. This shows that the pruning ratio settings do affect the NE results noticeably. Dividing a row into more chunks can potentially enable more fine-tuned pruning ratio settings and result in better NE, but at the same time may incur additional memory cost and access latency, which will be discussed quantitatively in Section 5.2.2. Setting appropriate pruning ratios and the number of chunks is a trade-off between NE, memory usage and training time.
Figure 9: Model quality vs. pruning ratio. (a) Model A. (b) Model B.
5.2.2 Overhead analysis
Memory overhead For a single-chunk FIITED model (i.e., K=1, same as AdaEmbed), the memory overhead can be estimated by $3/D$, where $D$ is the embedding dimension and 3 accounts for storing 1 chunk address, 1 utility value and 1 access frequency value per chunk. For K$>1$, the memory overhead of FIITED grows in a linear scale and becomes $3K/D$. During inference, both the utility value and the frequency value can be discarded, resulting in a $K/D$ memory overhead.
@@ 5.2.2 Overhead analysis
Performance overhead Figure 8 shows the average training/inference time per iteration of MD embeddings, AdaEmbed, ESAPN and FIITED with $K=2,4,8$, using Criteo Kaggle dataset with a batch size of 4096. The reported numbers are divided by the per-iteration time taken by an unpruned model for normalization purposes. As shown in the figure, during training, when $K=2$, FIITED is able to outperform MD embeddings and the unpruned DLRM, while seeing a small overhead compared to AdaEmbed (i.e., FIITED with $K=1$) . The performance overhead grows linearly to the number of chunks. FIITED is significantly faster than ESAPN. During inference, the overhead of FIITED remains smaller than MD embeddings and ESAPN, and is slightly larger than AdaEmbed.
6 Conclusion
@@
In this paper, we propose FIITED, an in-training embedding dimension optimization method that is able to directly cut down training memory footprint of DLRMs. Given a memory budget, FIITED can be plugged in directly into training without any need for prior knowledge of training data, pre-training or re-training. Embedding dimensions are adjusted during training in a fine-grained manner while changing data statistics are taken into consideration. Experiments on two industry models show that FIITED consistently achieves better NE than a state-of-the-art in-training embedding pruning method given the same average pruning ratio, and can prune much more than the baseline (65% vs. 50% for Model A, 50% vs. 30% for Model B) without affecting evaluation NE. On public datasets, FIITED is able to achieve an embedding size pruning ratio up to 93.75%-99.75% on three click-through-rate datasets and 77.5%-81.25% on one classification dataset, without significant accuracy loss.

### v2

@@ IV-A Methodology
Figure 10: (a) Average training/inference time per iteration of different models. (b) Ablation study result of overhead reduction design in FIITED. Both experiments on Criteo Kaggle dataset with batch size=4096.
IV-B Memory Saving Results
@@
Public models. For the public models, the criterion for model quality is prediction accuracy. According to  [34] and  [22], an accuracy loss larger than 0.1%-level is considered significant for click-through-rate prediction tasks. Table II shows the minimum equivalent number of entries that AdaEmbed, ESAPN, and FIITED are able to keep in the embedding table, while retaining the accuracy budget required by each task. On the three click-through-rate prediction datasets (i.e., Criteo Kaggle, Criteo Terabyte, and Avazu), the accuracy budget is set to less than 0.1% lower than the unpruned model, while on the MovieLens-20M dataset, we reported the minimum number of preserved entries to obtain a comparable accuracy to ESAPN. On all four datasets, FIITED can keep fewer embedding entries than the baselines while maintaining the model quality. Overall, FIITED is able to achieve 100×100$\times$100 × to 400×400$\times$400 × embedding table reduction on the three click-through-rate prediction datasets, yielding 1.5×1.5$\times$1.5 × to 4×4$\times$4 × improvement compared to AdaEmbed, and $\times2$ to 2.7×2.7$\times$2.7 × on the classification dataset, a 1.05×1.05$\times$1.05 × to 1.3×1.3$\times$1.3 × improvement compared to ESAPN. Figure 9 shows the validation accuracy comparison of FIITED, AdaEmbed and Mixed Dimension on the Criteo Kaggle dataset under different model size reduction ratios. FIITED consistently achieves better accuracy than the baselines, and the two pruning ratio selection approaches (manual and dynamical) yield similar results.
We then verify the performance of 2D pruning using the same public datasets mentioned above. The 2D pruning method is implemented upon the manual pruning ratio setting of FIITED. In Table III, we report the minimum preserving result using the 2D pruning method of FIITED. In the table, we showcased the original number of entries and dimension size used by the four public datasets, and the average number of preserved entries and preserved dimension size used by FIITED-2D to achieve the best pruning result while meeting the accuracy requirement on each dataset, with the corresponding overall reduction ratio. Compared with the reduction results of FIITED (Manual) in Table III, 2D pruning is able to achieve a higher overall reduction ratio on all four datasets, especially on Avazu and MovieLens-20M, where dimension size can be significantly reduced without harming accuracy. Overall, FIITED-2D can achieve 1.3×1.3$\times$1.3 × to $\times6$ improvements in embedding savings compared to FIITED (Manual), and 2×2$\times$2 × to $\times9$ improvements compared to AdaEmbed and ESAPN.
@@
Production models. The criterion for model quality is Normalized Entropy (NE) [8, 19] on the evaluation set; a lower NE indicates a better model. The effect of pruning is evaluated by NE loss, i.e., the percentage change in NE after model pruning compared to an unpruned model. A positive NE loss indicates worsened model quality, and an NE loss bigger than 0.02% is generally considered significant. The results are shown in Figure 9. For both Model A and Model B, FIITED can achieve better NE than AdaEmbed given the same average pruning ratio. For Model A, without incurring significant NE loss ($>$0.02%), FIITED can prune 65%-75% of the embeddings, while AdaEmbed can only prune 50%-65%. For Model B, FIITED can prune up to 50% of the embeddings, while AdaEmbed can prune 30%-40% of the embeddings without significant NE loss.
@@
For Model A, we also compare two different per-chunk pruning ratio settings with the same average ratio (75%). The “steeper” setting achieved a better NE (+0.02% VS. +0.03%) than the “flatter” setting. This shows that per-chunk pruning ratios do affect model quality noticeably. Dividing a row into more chunks can potentially enable more fine-tuned pruning ratio settings and result in better NE, but at the same time may incur additional memory cost and access latency (Section IV-C). Setting appropriate pruning ratios and the number of chunks is a trade-off between model quality, memory usage, and training time.
IV-C Overhead analysis
@@
End-to-end training overhead. Figure 10(a) shows the average training/inference time per iteration of MD embeddings, AdaEmbed, ESAPN, and FIITED with $K=2,4,8$, using Criteo Kaggle dataset with a batch size of 4096. The reported numbers are divided by the per-iteration time taken by an unpruned model for normalization purposes. As shown in the figure, during training, when $K=2,4$, FIITED is able to outperform MD embeddings and the unpruned DLRM, yielding a 1.41$\times$ and 1.13$\times$ speedup compared with the unpruned DLRM, while seeing a small overhead compared to AdaEmbed (i.e., FIITED with $K=1$) and being significantly faster than ESAPN. The performance overhead steadily grows to the number of chunks, remaining faster or negligibly slower than DLRM. During inference, the overhead of FIITED remains smaller than MD embeddings and ESAPN, and is slightly larger than AdaEmbed. Overall, the end-to-end training and inference performance of FIITED stays faster or comparable to the unpruned model (DLRM), and the overhead grows in a controllable scale to the number of chunks.
Figure 11: Memory overhead of FIITED on production models.
Figure 12: Pruning ratios used in the simulation when the number of chunks is 8.
Figure 13: Embedding accesses comparison (left: number of chunk accesses; right: total amount of accessed memory, including repeated accesses). Normalization is performed by dividing values by the number of embedding accesses in an unpruned model.
Memory overhead. For a single-chunk FIITED model (i.e., $K=1$, same as AdaEmbed), the memory overhead can be estimated by $3/D$, where $D$ is the embedding dimension and 3 accounts for storing 1 chunk address, 1 utility value and 1 access frequency value per chunk. For $K>1$, the memory overhead of FIITED grows in a linear scale and becomes $3K/D$. During inference, both the utility value and the frequency value can be discarded, resulting in a $K/D$ memory overhead. To further verify the memory overhead of FIITED, we ran a simulation of the VHPI embedding store system on the Kaggle dataset, which will be discussed in detail in Section IV-D. The memory overhead of FIITED in the simulation is shown in Figure 12, where the memory overhead scales approximately linearly to the number of chunks.
Figure 14: Embedding evictions comparison (left: number of chunk evictions; right: total amount of evicted memory). Normalization is performed by dividing values by the number of unique embeddings in an unpruned model.
Figure 15: Embedding allocations comparison (left: number of chunk allocations; right: total amount of allocated memory, including re-allocations). Normalization is performed by dividing values by the number of unique embeddings in an unpruned model.
IV-D Ablation Study
TABLE IV: Comparison of existing EDS methods with FIITED.
EDS Comparison Pre-training Post-training In-training (except FIITED) FIITED
Require prior knowledge of training data Yes No No No
Reduce training memory usage Yes No No Yes
Adapt to dynamic changes in data traits No No Yes Yes
Impact of training overhead optimization. To verify the effectiveness of our overhead optimization method proposed in Section III-D, we further conduct ablation studies on FIITED regarding the optimization effect of parallelization and pipeline overlapping. Figure 10(b) shows the average end-to-end per-iteration training time of FIITED with $K=2,4,8$ under the same settings as described above, compared with the system implementations with only parallelization or overlapping. All reported numbers are normalized using the time taken from the unpruned DLRM model, using the same approach as the experiment results described above. The result clearly shows the efficacy of the system design of FIITED, representing significant performance reduction compared with the system designs with only parallelization or overlapping. In the experiment of public datasets, we only used 1 GPU in order to simulate the scenario with restricted computation resources. Hence, parallelization and pipeline overlapping are only done using multi-threading and multiprocessing on a single device. On a larger distributed system, the performance overhead of FIITED can be further reduced by distributing the retrieval and utility update of different chunks on different devices, following the same scheme as shown in this section.
Impact of VHPI embedding table. In order to further analyze the overhead of FIITED, a simulation of the VHPI embedding store system is implemented in Python and run on 7 days of data in the open-source Kaggle dataset. All the VHPI embedding store operations are simulated and a full-size VHPI hash table is maintained. The VHPI embedding table is simplified in the simulation and does not contain actual embeddings. Training is not done during the simulation.
To specify the pruning ratio for each embedding chunk, a power law function $f(x)=cx^{a}$ at discrete points $x=(i+0.5)/K,i=0,1,2,...,K-1$ in the interval [0,1] is fitted to satisfy the desired average pruning ratio. We choose 2 settings for the power law parameter $a$: $a=0.5$ for a concave power function and $a=1$ for a linear power function, and compute the normalization coefficient $c$ during curve fitting. If a chunk’s pruning ratio is too big (bigger than 95%) after curve fitting, it is reduced to 95%, and the curve fitting process is re-run for the rest of the chunks. The average pruning ratio is set to 50%, 70%, and 90% in the experiments. The pruning ratios used in the simulation when $K=8$ are plotted in Figure 12. At 90% average pruning ratio, due to limiting the per chunk pruning ratio to a maximum of 95%, the two power-law settings ($a=0.5$ and $a=1$) yield the same pruning ratios.
We plot the statistics of embedding accesses, evictions, and allocations in Figures 13, 14 and 15. The left-hand sides in the figures show the total number of operations performed at the chunk level, while the right-hand sides display the averaged number of operations performed at the embedding level. The right-hand sides can be derived from the left-hand sides by dividing each value by the corresponding number of chunks. The values are normalized according to statistics of an unpruned model.
V Related Work
V-A Embedding Dimension Search (EDS)
Existing EDS approaches can mainly be classified into three categories: (1) Pre-training, where embedding dimensions are decided before the actual training according to information extracted from the training dataset or a lightweight pre-training process  [20, 15, 5]. Because the selected dimensions do not change during training, pre-training EDS misses opportunities to take advantage of changing data characteristics during training. Pre-training EDS also relies heavily on prior knowledge of the training data, which is not required by FIITED. (2) Post-training, where dimension pruning is performed after training. Additional information may be collected during training to aid pruning, and re-training is sometimes needed to boost the model performance  [1, 32, 10]. Post-training EDS cannot reduce training memory footprint. (3) In-training, where the pruning decisions are made during training. Existing works  [14, 26, 33] add additional network structures on top of the original recommender model, e.g., new DNN layers, which help select embedding dimensions during training, but the majority of them still need to store embeddings at the maximum dimension at the time of training and thus do not reduce training memory usage. One exception is ESAPN [14], which stores embeddings at their current lengths during training by abandoning the old embeddings in memory whenever the embedding dimensions are increased. However, without providing a system design, it is unclear how ESAPN can utilize the small pieces of free memory. ESAPN also does not provide control over how much memory is used during training, while FIITED can reduce memory usage to an arbitrary desired amount. In addition, ESAPN trains one policy network per sparse feature and introduces significant training time overhead, while FIITED has negligible runtime overhead.
V-B Embedding Compression
To decrease the size of embedding tables, one can: (1) reduce the size of each value in the table; (2) reduce the number of rows, i.e., the hash size; (3) reduce the number of columns, i.e., the embedding dimension; or (4) deconstruct the embedding table by replacing the traditional 2D array storage format with novel designs. (1) is commonly achieved by quantization  [30, 7, 27]. (2) involves designing novel hashing methods to reduce the number of rows while maintaining the model quality  [29, 25, 3]. (3) has been discussed previously in Section V-A and includes EDS methods performed prior to, during or after training. As shown in Table IV, unlike existing EDS methods, FIITED is able to reduce the training memory footprint while adjusting embedding dimensions based on dynamic changes in data characteristics. For (4), novel embedding storage designs in existing research include constructing embeddings from two separate tables  [18], via multi-layer embeddings  [4], and by applying transformation matrices to a small set of anchor embeddings  [13].
VI Conclusion
@@
In this paper, we propose FIITED, an in-training embedding dimension optimization method that is able to directly cut down the training memory footprint of DLRMs. Given a memory budget, FIITED can be plugged directly into training without any need for prior knowledge of training data, pre-training, or re-training. Embedding dimensions are adjusted during training in a fine-grained manner while changing data statistics are taken into consideration. Experiments on two industry models show that FIITED consistently achieves better NE than a state-of-the-art in-training embedding pruning method given the same average pruning ratio, and can prune much more than the baseline (65% vs. 50% for Model A, 50% vs. 30% for Model B) without affecting evaluation NE. On public datasets, FIITED can reduce the embedding size by 100$\times$ to 800$\times$ on three click-through-rate datasets and $\times2$ to 17.8×17.8$\times$17.8 × on one classification dataset, without significant accuracy loss.